# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ I-E Paper Organization
The remainder of this paper is organized as below. Section II discusses the working principles of single LM agents, while Section III presents the cooperation paradigms for connected LM agents. Section IV and Section V introduce the taxonomy of security and privacy threats to LM agents, respectively, along with state-of-the-art countermeasures. Section VI outlines open research issues and future directions in the field of LM agents. Finally, conclusions are drawn in Section VII. Fig. 4 depicts the organization structure of this survey.
@@
and agents to effectively communicate with LLM-enabled agents, such as API syntax and semantics.
@@
The integration of LLMs with existing educational systems ensures that LLMs can seamlessly interact with AIS
while addressing issues of bias, transparency, and accountability within educational environments.
II Large Model Agents: Working Principles
In this section, we first introduce existing standards of LM agents. Then, we discuss the general architecture of connected LM agents including key components, communication modes, key characteristics, and enabling technologies. Next, we introduce typical prototypes and discuss modern applications of LM agents.
@@ II-A Standards of LM Agents
1) The IEEE SA - P3394 standard [^1] https://standards.ieee.org/ieee/3394/11377/, launched in 2023, defines natural language interfaces to facilitate communication between LLM applications, agents, and human users. This standard establishes a set of protocols and guidelines that enable applications and agents to effectively communicate with LLM-enabled agents. These protocols and guidelines include, but are not limited to, API syntax and semantics, voice and text formats, conversation flow, prompt engineering integration, LLM thought chain integration, as well as API endpoint configuration, authentication, and authorization for LLM plugins. The standard is expected to advance technological interoperability, promote AI industry development, enhance the practicality and efficiency of LMs, and improve AI agent functionality and user experience.
@@
2) The IEEE SA - P3428 standard [^2] https://standards.ieee.org/ieee/3428/11489/, launched in 2023, aims to develop standards for LLM agents in educational applications. The primary goal is to ensure the interoperability of LLM agents across both open-source and proprietary systems. Key areas of focus include the integration of LLMs with existing educational systems and addressing technical and ethical challenges. This includes ensuring that LLMs can seamlessly interact with other AI components, such as Adaptive Instructional Systems (AIS), while also addressing issues of bias, transparency, and accountability within educational contexts. The standard is intended to support the widespread and effective application of LLMs in the field of education, thereby enabling more personalized, efficient, and ethically sound AI-driven educational experiences.
Figure 6: General Architecture of Internet of LM agents in bridging the human, physical, and cyber worlds. The LM agent has five constructing modules: planning, action, memory, interaction, and security modules. The engine of connected LM agents is empowered by a combination of five cutting-edge technologies: large foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration. LM agents interact with humans through Human-Machine Interaction (HMI) technologies such as NLP, with the assistance of handheld devices and wearables, to understand human intentions, desires, and beliefs. LM agent can synchronize data and statuses between the physical body and the digital brain through digital twin technologies, as well as perceive and act upon the surrounding virtual/real environment. LM agents can be interconnected in the cyber space through efficient cloud-edge networking for efficient data and knowledge sharing to promote multi-agent collaboration.
II-B Architecture of Connected LM Agents
II-B1 Operating System (OS) of LM Agents
@@
According to [45, 46], the OS architecture of LM agents consists of three layers: application, kernel, and hardware.
@@
  • The application layer hosts agent applications (e.g., travel, coding, and robot agents) and offers a SDK that abstracts system calls, simplifying agent development.
@@
  • The kernel layer includes the ordinary OS kernel and an additional LM agent kernel, with a focus on without altering the original OS structure. Key modules in the LM agent kernel [45, 46] include the agent scheduler for task planning and prioritization, context manager for LM status management, memory manager for short-term data, storage manager for long-term data retention, tool manager for external API interactions, and access manager for privacy controls.
@@
II-B2 Constructing Modules of LM Agents
According to [8, 1, 9, 10], there are generally five constructing modules of LM agents: planning, action, memory, interaction, and security modules (details in Sect. II-C). These modules together enable LM agents to perceive, plan, act, learn, and interact efficiently and securely in complex and dynamic environments.
  • Empowered by LMs, the planning module produces strategies and action plans with the help of the memory module, enabling informed decision-making [7, 10].
  • The action module executes these embodied actions, adapting actions based on real-time environmental feedback to ensure contextually appropriate responses [9, 42].
  • The memory module serves as a repository of accumulated knowledge (e.g., past experiences and external knowledge), facilitating continuous learning and improvement [10].
  • The interaction module enables effective communication and collaboration with humans, other agents, and environment.
  • The security module is integrated throughout LM agents’ operations, ensuring active protection against threats and maintaining integrity and confidentiality of data and processes.
II-B3 Engine of LM Agents
The engine of LM agents is powered by a combination of cutting-edge technologies including large foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration (details in Sect. II-D).
  • Large foundation model such as GPT-4 and DALL-E 2 serves as the brain of an LM agent, which enables high-level pattern recognition, advanced reasoning, and intelligent decision-making, providing the cognitive capabilities of LM agents [6, 7].
  • Knowledge-related technologies enhance LM agents by incorporating Knowledge Graphs (KGs), knowledge bases, and RAG systems, allowing agents to access, utilize, and manage vast external knowledge sources, ensuring informed and contextually relevant actions [47].
  • HMI technologies enable seamless interaction between humans and agents through NLP, multimodal interfaces, and Augmented/Virtual/Mixed Reality (AR/VR/MR), facilitating dynamic and adaptive interactions [48].
  • Digital twin technologies allows efficient and seamless synchronization of data and statuses between the physical body and the digital brain of an LM agent through intra-agent communications [49].
  • Multi-agent collaboration technologies empower LM agents to work together efficiently, sharing data, resources, and tasks to tackle complex problems by developing cooperation, competition, and coopetition strategies through inter-agent communications [28].
II-B4 Communication Mode of LM Agents
Every LM agent consists of two parts: (i) the LM-empowered brain located in the cloud, edge servers, or end devices and (ii) the corresponding physical body such as autonomous vehicle. Every LM agent can actively interact with other LM agents, the virtual/real environment, and humans. For connected LM agents, there exist two typical communication modes: intra-agent communications for seamless data/knowledge synchronization between brain and physical body within an LM agent, and inter-agent communications for efficient coordination between LM agents. Table III summarizes the comparison of the two communication modes.
  • Intra-agent communications refer to the internal data/knowledge exchange within a single LM agent. This type of communication ensures that different components of the LM agent, including planning, action, memory, interaction, and security modules, work in harmony. For example, an LM agent collects multimodal sensory data through its physical body, which then communicates the interpreted information to the LM-empowered brain. The planning module in the brain formulates a response or action plan, which is then executed by the action module. This seamless flow of information is critical for maintaining the LM agent’s functionality, coherence, and responsiveness in real-time and dynamic scenarios.
  • Inter-agent communications involve information and knowledge exchange between multiple LM agents. It enables collaborative task allocation, resource sharing, and coordinated actions among agents to foster collective intelligence. For example, in a smart city application, various LM agents managing traffic lights, public transportation, and emergency services share real-time data to optimize urban mobility and safety. Effective inter-agent communications rely on standardized protocols to ensure compatibility and interoperability, facilitating efficient and synchronized operations across the network of LM agents.
TABLE III: A Summary of Intra-Agent And Inter-Agent Communications for Connected LM Agents
Intra-agent Comm. Inter-agent Comm.
[…]
Connection Type Within a single LM agent Among multiple LM agents
Support Two-way Communication
Support Multimodal Interaction
Support Semantic Communication
Typical Communication Environment Wireless Wired
II-B5 Information Flow Between Human World and LM Agents
Humans interact with LM agents through natural language, mobile smart devices, and wearable technology, enabling LM agents to comprehend human instructions and address real-world issues effectively. LM agents, in turn, acquire new knowledge and data from human inputs, which aids in their continuous improvement and learning. This ongoing process of updating and optimizing their models allows LM agents to provide increasingly accurate and useful information. In AR and VR environments, LM agents can work collaboratively with human users in virtual settings, such as architectural design, for enhanced overall efficiency and creativity [50].
II-B6 Information Flow Between Physical World and LM Agents
Empowered by digital twin technologies, LM agents are allowed to synchronize data and statuses between their physical bodies and their digital brains, creating a seamless interaction loop. LM agents can also monitor and act upon real-time inputs from their environments. This bidirectional synchronization allows LM agents to perceive and respond to their surrounding environments—whether virtual or real—with a high degree of precision and responsiveness, thus bridging the gap between the digital and physical realms. By continuously learning from environmental feedbacks, LM agents can accumulate knowledge and develop an understanding of physical laws, which empowers them to solve complex real-world problems. This iterative learning process ensures that LM agents not only react to immediate stimuli but also refine their embodied actions over time, achieving more sophisticated and effective solutions.
II-B7 Information Flow Between Cyber World and LM Agents
In the cyber world, LM agents are interconnected into the Internet of LM agents through efficient cloud-edge networking, facilitating seamless data and knowledge sharing that promotes multi-agent collaboration. By deploying LMs across both cloud and edge infrastructures, it allows LM agents to leverage the strengths of both cloud and edge computing for optimized performance and responsiveness [51]. The cloud provides substantial computational power and storage, enabling the processing of vast amounts of data and the training of sophisticated models. Meanwhile, the edge offers real-time data processing capabilities closer to the source, reducing latency and ensuring timely decision-making. In the Internet of LM agents, LM agents can collaboratively share data, knowledge, and learned experiences with others in real-time, creating a robust and adaptive network of intelligence across multiple domains. For example, in a smart city, embodied LM agents in various locations can work together to optimize traffic flow, manage energy resources, and enhance public safety by sharing real-time data and coordinating their actions.
@@ II-B7 Information Flow Between Cyber World and LM Agents
II-C Key Components of LM Agent
As depicted in Fig. 7, it generally contains five key modules to construct a connected LM agent [8, 1, 9, 10].
@@ II-C1 Planning Module
The planning module serves as the core of an LM agent [7, 10]. It utilizes advanced reasoning techniques to enable LM agents to devise efficient and effective solutions to complex problems. The working modes of the planning module include the following types.
  • Feedback-free planning: The planning module enables LM agents to understand the complex problems and find reliable solutions by breaking them down into necessary steps or manageable sub-tasks [7, 14]. For example, CoT [11] is a popular sequential reasoning approach where each thought builds directly on the previous one. It represents the step-by-step logical thinking and can enhance the generation of coherent and contextually relevant responses. ToT [12] organizes reasoning as a tree-like structure, exploring multiple paths simultaneously. In ToT, Each node represents a partial solution, allowing the model to branch and backtrack to find the optimal answer. Graph of Thought (GoT) [52] models reasoning using an arbitrary graph structure, allowing more flexible information flow. GoT captures complex relationships between thoughts, enhancing the model’s problem-solving capabilities. AVIS [53] further refines the tree search process for visual QA tasks using a human-defined transition graph and enhances decision-making through a dynamic prompt manager.
  • Feedback-enhanced planning: To make effective long-term planning in complex tasks, it is necessary to iteratively reflect on and refine execution plans based on past actions and observations [39]. The goal is to correct past errors and improve final outcomes. For example, ReAct [54] combines reasoning and acting by prompting LLMs to generate reasoning traces and actions simultaneously. This dual approach allows the LLM to create, monitor, and adjust action plans, while task-specific actions enhance interaction with external sources, improving response accuracy and reliability. Reflexion [55] converts environmental feedback into self-reflection and enhances ReAct by enabling LLM agents to learn from past errors and iteratively optimize behaviors. Reflexion features an actor that produces actions and text via models (e.g., CoT and ReAct) enhanced by memory, an evaluator that scores outputs using task-specific reward functions, and self-reflection that generates verbal feedback to improve the actor.
@@
  • Grounded planning: Executing plans in real-world environments (e.g., Minecraft) requires precise, multi-step reasoning. VOYAGER [50], the first LLM-powered agent in Minecraft, utilizes in-context lifelong learning to adapt and generalize skills to new tasks and worlds. VOYAGER include an automatic curriculum for exploration, a skill library of executable code for complex behaviors, and an iterative prompting mechanism that refines programs based on feedback. Wang et al. [57] further propose an interactive describe-explain-plan-select (DEPS) planning approach that improves LLM-generated plans by integrating execution descriptions, self-explanations, and a goal selector that ranks sub-goals to refine planning. Additionally, Song et al. [7] present a grounded re-planning algorithm which dynamically updates high-level plans during task based on environmental perceptions, triggering re-planning when actions fail or after a specified time.
@@ II-C2 Memory Module
The memory module is integral to LM agent’s ability to learn and adapt over time [39]. It maintains an internal memory that accumulates knowledge from past interactions, thoughts, actions, observations, and experiences with users, other agents, and the environments. The stored information guides future decisions and actions, allowing the agent to continuously refine its knowledge and skills. This module ensures that the agent can remember and apply past lessons to new situations, thereby improving its long-term performance and adaptability [10]. There are various memory formats such as natural language, embedded vectors, databases, and structured lists. Additionally, RAG technologies [15] are employed to access external knowledge sources, further enhancing the accuracy and relevance of LM agent’s planning capabilities. In the literature [39, 10], memory can be divided into the following three types.
@@
  • Hybrid memory combines short-term and long-term memory to enhance an agent’s understanding of the current context and leverage past experiences for better long-term reasoning. Liu et al. [59] propose the RAISE architecture to enhance ReAct for conversational AI agents by integrating a dual-component memory system, where Scratchpad captures recent interactions as short-term memory; while the retrieval module acts as long-term memory to access relevant examples. HIAGENT [60] employs cross-trial and in-trial memory, where cross-trial memory stores historical trajectories and in-trial memory captures current trials. Instead of retaining all action-observation pairs, HIAGENT uses subgoals as memory chunks to save memory, each containing summarized observations. LLM generates subgoals, executes actions to achieve them, and updates the working memory by summarizing and replacing completed subgoals with relevant information.
@@ II-C3 Action Module
  • Embodied actions. The action module empowers LM agents to perform contextually appropriate embodied actions and adapt to environmental changes, facilitating interaction with and adjustment to physical surroundings [25, 21]. As LLM-generated action plans are often not directly executable in interactive environments, Huang et al. [25] propose refining LLM-generated plans for embodied agents by conditioning on demonstrations and semantically translating them into admissible actions. Evaluations in the VirtualHome environment show significant improvements in executability, ranging from 18% to 79% over the baseline LLM. Besides, SayCan [21] enables embodied agents such as robots to follow high-level instructions by leveraging LLM knowledge in physically-grounded tasks, where LLM (i.e., Say) suggests useful actions; while learned affordance functions (i.e., Can) assess feasibility. SayCan’s effectiveness is demonstrated through 101 zero-shot real-world robotic tasks in a kitchen setting. PaLM-E [61] is a versatile multimodal language model for embodied reasoning, visual-language, and language tasks. It integrates continuous sensor inputs, e.g., images and state estimates, into the same embedding space as language tokens, allowing for grounded inferences in real-world sequential decision-making.
  • Learning to use & make tools. By leveraging various tools (e.g., search engines and external APIs) [62], LM agent can gather valuable information to handle assigned complex tasks. For example, AutoGPT integrates LLMs with predetermined tools such as web and file browsing. InteRecAgent [63] integrates LLMs as the brain and recommender models as tools, using querying, retrieval, and ranking tools to handle complex user inquiries. Beyond using existing tools, LM agents can also develop new tools to enhance task efficiency [9]. To optimize tool selection with a large toolset, ReInvoke [64] introduces an unsupervised tool retrieval method featuring a query generator to enrich tool documents in offline indexing and an intent extractor to identify tool-related intents from user queries in online inference, followed by a multi-view similarity ranking strategy to identify the most relevant tools.

### v2

@@ I-E Paper Organization
The remainder of this paper is organized as below. Section II gives an overview of single LM agents, while Section III presents the fundamentals for networking LM agents. Section IV and Section V introduce the taxonomy of security and privacy threats to LM agents, respectively, along with state-of-the-art countermeasures. Section VI outlines open research issues and future directions in the field of LM agents. Finally, conclusions are drawn in Section VII. Fig. 4 depicts the organization structure of this survey.
@@
and agents to effectively communicate with LLM agents, such as API syntax and semantics.
@@
The integration of LLMs with existing education systems to ensure that LLMs can seamlessly interact with
AIS while addressing issues of bias, transparency, and accountability in educational environments.
II An Overview of Large Model Agents
In this section, we first introduce existing standards of LM agents. Then, we discuss the operating system (OS) and constructing modules of LM agents. Next, we introduce typical prototypes and discuss modern applications of LM agents.
@@ II-A Standards of LM Agents
1) The IEEE SA - P3394 standard [^1] https://standards.ieee.org/ieee/3394/11377/, launched in 2023, defines natural language interfaces to facilitate communication between LLM applications, agents, and human users. This standard establishes a set of protocols and guidelines that enable applications and agents to effectively communicate with LLM agents. These protocols and guidelines include, but are not limited to, API syntax and semantics, voice and text formats, conversation flow, prompt engineering integration, LLM thought chain integration, as well as API endpoint configuration, authentication, and authorization for LLM plugins. The standard is expected to advance technological interoperability, promote AI industry development, enhance the practicality and efficiency of LMs, and improve AI agents’ functionality and user experience.
@@
2) The IEEE SA - P3428 standard [^2] https://standards.ieee.org/ieee/3428/11489/, launched in 2023, aims to develop standards for LLM agents in educational applications. The primary goal is to ensure the interoperability of LLM agents across both open-source and proprietary systems. Key areas of focus include the integration of LLMs with existing educational systems and addressing technical and ethical challenges. This includes ensuring that LLMs can seamlessly interact with other AI components, such as adaptive instructional systems (AIS), while addressing issues of bias, transparency, and accountability within educational contexts. The standard is intended to support the widespread and effective application of LLMs in the field of education, thereby enabling more personalized, efficient, and ethically sound AI-driven educational experiences.
II-B Operating System (OS) of LM Agents
@@
According to [47, 48], the OS architecture of LM agents consists of three layers: application, kernel, and hardware, as shown in Fig. 5.
@@
  • The application layer hosts agent applications (e.g., travel, coding, and robot agents) and offers an SDK that abstracts system calls, simplifying agent development. Typical agent developing frameworks include LangChain, AutoGen, Dify, and LlamaIndex.
@@
  • The kernel layer includes the ordinary OS kernel (to process LM-related queries such as file operation and network request) and an additional LM agent kernel (to process LM-related queries such as reasoning, planning, and tool parsing), with a focus on without altering the original OS structure. Key modules in the LM agent kernel [47, 48] include the agent scheduler for task planning and prioritization, context manager for LM status management, memory manager for short-term data, storage manager for long-term data retention, tool manager for external API interactions, and access manager for privacy controls.
@@
Figure 6: Workflow of the five constructing modules of LM agents in interacting with the real/virtual environment.
II-C Constructing Modules of LM Agents
According to [9, 2, 10, 11], there are generally five constructing modules of LM agents: planning, action, memory, interaction, and security. As depicted in Fig. 6, these modules together enable LM agents to perceive, plan, act, learn, and interact efficiently and securely in complex and dynamic environments. Particularly, empowered by LMs, the planning module produces strategies and action plans with the help of the memory module, enabling informed decision-making. The action module executes these embodied actions, adapting actions based on real-time environmental feedback to ensure contextually appropriate responses. The memory module serves as a repository of accumulated knowledge (e.g., past experiences and external knowledge), facilitating continuous learning and improvement. The interaction module enables effective communication and collaboration with humans, other agents, and environment. The security module is integrated throughout LM agents’ operations, ensuring active protection against threats and maintaining integrity and confidentiality of data and processes.
@@ II-C Constructing Modules of LM Agents
Figure 8: Illustration of typical feedback-free planning methods of LMs: (a) basic I/O; (b) chain-of-thought (CoT) [49]; (c) tree-of-thought (ToT) [50]; (d) graph-of-thought (GoT) [51].
Figure 9: Illustration of ReAct [52]: a typical feedback-enhanced planning methods of LMs.
Figure 10: Illustration of typical feedback-enhanced planning methods of LMs: (a) Reflexion [53]; (b) ground planning [54].
@@ II-C1 Planning Module
The planning module serves as the core of an LM agent [8, 11]. It utilizes advanced reasoning techniques to devise effective solutions fo agents to complex problems. It includes the following working modes.
  • Feedback-free planning: The planning module enables LM agents to understand the complex problems and find reliable solutions by breaking them down into necessary steps or manageable sub-tasks [8, 12], as shown in Fig. 8. For instance, chain-of-thought (CoT) [49] is a popular sequential reasoning approach where each thought builds directly on the previous one. It represents the step-by-step logical thinking and can enhance the generation of coherent and contextually relevant responses. Tree-of-thought (ToT) [50] organizes reasoning as a tree-like structure, exploring multiple paths simultaneously. In ToT, each node represents a partial solution, allowing the model to branch and backtrack to find the optimal answer. Graph-of-thought (GoT) [51] models reasoning using an arbitrary graph structure, allowing more flexible information flow. GoT captures complex relationships between thoughts, enhancing the model’s problem-solving capabilities. AVIS [55] further refines the tree search process for visual QA tasks using a human-defined transition graph and enhances decision-making through a dynamic prompt manager.
  • Feedback-enhanced planning: To make effective long-term planning in complex tasks, it is necessary to iteratively reflect on and refine execution plans based on past actions and observations [41]. The goal is to correct past errors and improve final outcomes. For instance, as shown in Fig. 9, ReAct [52] combines reasoning and acting by prompting LLMs to generate reasoning traces (i.e., thoughts) and actions simultaneously. This dual approach allows the LLM to create, monitor, and adjust action plans, while task-specific actions enhance interaction with external sources, improving response accuracy and reliability. As shown in Fig. 10(a), Reflexion [53] converts environmental feedback into self-reflection and enhances ReAct by enabling LLM agents to learn from past errors and iteratively optimize behaviors. Reflexion features an actor that produces actions and text via models (e.g., CoT and ReAct) enhanced by memory, an evaluator that scores outputs using task-specific reward functions, and self-reflection that generates verbal feedback to improve the actor.
@@
  • Grounded planning: Executing plans in real or simulated world environments (e.g., Minecraft) requires precise, multi-step reasoning. As shown in Fig. 10(b), VOYAGER [54], the first LLM-powered agent in Minecraft, utilizes in-context lifelong learning to adapt and generalize skills to new tasks and worlds. VOYAGER include an automatic curriculum for exploration, a skill library of executable code for complex behaviors, and an iterative prompting mechanism that refines programs based on feedback. Wang et al. [57] further propose an interactive describe-explain-plan-select (DEPS) planning approach that improves LLM-generated plans by integrating execution descriptions, self-explanations, and a goal selector that ranks sub-goals to refine planning.
@@ II-C2 Memory Module
The memory module is integral to LM agent’s ability to learn and adapt over time [41]. It maintains an internal memory that accumulates knowledge from past interactions, thoughts, actions, observations, and experiences with users, other agents, and the environments. The stored information guides future decisions and actions, allowing the agent to continuously refine its knowledge and skills. This module ensures that the agent can remember and apply past lessons to new situations, thereby improving its long-term performance and adaptability [11]. There are various memory formats such as embedded vectors, databases, and structured lists. Additionally, retrieval-augmented generation (RAG) technologies [58] are employed to access external knowledge sources, further enhancing the accuracy and relevance of LM agent’s planning capabilities. In the literature [41, 11], memory can be divided into the following three types.
@@
  • Hybrid memory combines short-term and long-term memory to enhance an agent’s understanding of the current context and leverage past experiences for better long-term reasoning. Liu et al. [60] propose the RAISE architecture to enhance ReAct for conversational AI agents by integrating a dual-component memory system, where Scratchpad captures recent interactions as short-term memory; while the retrieval module acts as long-term memory to access relevant examples.
@@ II-C3 Action Module
  • Embodied actions. The action module empowers LM agents to perform contextually appropriate embodied actions and adapt to environmental changes, facilitating interaction with and adjustment to physical surroundings [61, 15]. As LLM-generated action plans are often not directly executable in interactive environments, Huang et al. [61] propose refining LLM-generated plans for embodied agents by conditioning on demonstrations and semantically translating them into admissible actions. Evaluations in the VirtualHome environment show significant improvements in executability, ranging from 18% to 79% over the baseline LLM. Besides, SayCan [15] enables embodied agents such as robots to follow high-level instructions by leveraging LLM knowledge in physically grounded tasks, where LLM (i.e., Say) suggests useful actions; while learned affordance functions (i.e., Can) assess feasibility. SayCan’s effectiveness is demonstrated through 101 zero-shot real-world robotic tasks in a kitchen setting. PaLM-E [62] is a versatile multimodal language model for embodied reasoning, visual-language, and language tasks. It integrates continuous sensor inputs, e.g., images and state estimates, into the same embedding space as language tokens, allowing for grounded inferences in real-world sequential decision-making.
  • Learning to use & make tools. By leveraging various tools (e.g., search engines and external APIs) [63], LM agents can gather valuable information to handle assigned complex tasks. For instance, AutoGPT integrates LLMs with predetermined tools such as web and file browsing. InteRecAgent [17] integrates LLMs as the brain and recommender models as tools, using querying, retrieval, and ranking tools to handle complex user inquiries. Beyond using existing tools, LM agents can also develop new tools to enhance task efficiency [10]. To optimize tool selection with a large toolset, ReInvoke [14] introduces an unsupervised tool retrieval method featuring a query generator to enrich tool documents in offline indexing and an intent extractor to identify tool-related intents from user queries in online inference, followed by a multi-view similarity ranking strategy to identify the most relevant tools.